OpenStudy (anonymous):

Find the additive and multiplicative inverse of -0.05.

OpenStudy (unklerhaukus):

if \(a'\) is the additive inverse of \(a\) iff \[a+(a')=0\] if \(b'\) is the multiplicative inverse of \(b\) iff \[b\times (b')=1\]

OpenStudy (unklerhaukus):

\[-0.05+a'=0\]\[\color{grey}{a'=}\] \[-0.05\times b'=1\]\[\color{grey}{b'=}\]